OpenStudy (anonymous):

Let f(x)=x^2+2x-15 determine the values for x for which f(x)= -16

OpenStudy (ajprincess):

since f(x)=-16 x^2+2x-15=-16 x^2+2x-15+16=0 x^2+2x+1=0 Can u find x nw?

OpenStudy (anonymous):

thanks :)

OpenStudy (ajprincess):

yw:)